OpenStudy (anonymous):

number of Mg(OH)2 formula units in 7.40 moles of Mg(OH)2

sam (.sam.):

7.40 moles of Mg(OH)2 Using the ratio, Multiply by the Avogadro number \[7.40 \cancel{\text{ moles of } Mg(OH)_2} \times \frac{6.02 \times 10^{23}\text{ atoms of } Mg(OH)_2}{1\cancel{\text{ mole of } Mg(OH)_2}}=4.45 \times 10^{24} \text{ atoms of } Mg(OH)_2\]

OpenStudy (anonymous):

Thanks. :)